OpenStudy (anonymous):

3x – 2(3x – 2) = 2(4 – x) + 3

OpenStudy (anonymous):

<=> 3x - 6x + 4 = 8 - 2x + 3 <=> 3x - 6x + 2x = 8 + 3 - 4 <=> -x = 7 <=> x = -7